OpenStudy (toxicsugar22):

hi can someone give me some elimatnition and substitution method to solve

OpenStudy (anonymous):

if u give a question i can explain

OpenStudy (anonymous):

or u want us to giv u questions??

OpenStudy (toxicsugar22):

yes i would prefer thatt

OpenStudy (toxicsugar22):

because i have a test tommorow and i want to do good so will u help me preper for that

OpenStudy (anonymous):

shud i give the two equations or should i give you a story problem where u hv to form the equations yourself

OpenStudy (toxicsugar22):

two equations please

OpenStudy (anonymous):

use elimination 9x - 4y = 2000 7x - 3y = 2000

OpenStudy (anonymous):

use substitution 3x 5y --- - ----- = -2 2 3 x y 13 ---- + ---- = ---- 3 2 6

OpenStudy (anonymous):

Hey toxicsugar, are u still there????? Shall i wait for yr reply??

OpenStudy (toxicsugar22):

im here srry

OpenStudy (toxicsugar22):

my computer froze

OpenStudy (anonymous):

working on the problems??

OpenStudy (toxicsugar22):

yes

OpenStudy (anonymous):

ok

OpenStudy (toxicsugar22):

wait can u help me on the first one i dont seem to get it

OpenStudy (toxicsugar22):

hakirat can u help me please

OpenStudy (anonymous):

ok it will be solved as follows: multiply first equation by 3 and second equation by 4 to get 27x -12y = 6000 -------(A) 28x - 12y = 8000 --------(B) Now subtract (B) from (A) to get -x = - 2000 so x = 2000 Now put x = 2000 in the given first equation to get value of y 9(2000) - 4y = 2000 -4y = 2000 - 18000 -4y = -16000 y = -16000/ -4 y = 4000

OpenStudy (toxicsugar22):

thaks

OpenStudy (anonymous):

were u able to solve the second one???

OpenStudy (toxicsugar22):

no not really

OpenStudy (anonymous):

I wish i cud answer now but it is past midnight in India and I hv to go to sleep... SORRY :}

OpenStudy (toxicsugar22):

oh your indian me to

OpenStudy (toxicsugar22):

which part of india do u live in

OpenStudy (anonymous):

I live in Delhi, the capital....